Use the Laplace transform to solve the given initial value problem.

y^(4)−256y=0; y(0)=43, y′(0)=104, y′′(0)=432, y′′′ (0)=1,024


Enclose arguments of functions in parentheses. For example, sin(2x).

y(t)= _______________
1
Expert's answer
2020-07-29T15:01:06-0400

"y(t)\to Y(p)"

"y^{(4)}(t)\to p^4Y(p)-p^3y(0)-p^2y'(0)-py''(0)-y^{(3)}(0)"

For our case:

"y^{(4)}(t)\to p^4Y(p)-43p^3-104p^2-432p-1024"

Then:

"p^4Y(p)-43p^3-104p^2-432p-1024-256Y(p)=0"

"(p^4-256)Y(p)=43p^3+104p^2+432p+1024"

"Y(p)=\frac{43p^3+104p^2+432p+1024}{(p^2+16)(p+4)(p-4)}"

"\frac{A}{p+4}+\frac{B}{p-4}+\frac{Cp+D}{p^2+16}=\frac{43p^3+104p^2+432p+1024}{(p^2+16)(p+4)(p-4)}"

"A(p-4)(p^2+16)+B(p+4)(p^2+16)+(Cp+D)(p+4)(p-4)="

"=43p^3+104p^2+432p+1024"

"A(p^3-4p^2+16p-64)+B(p^3+4p^2+16p+64)+Cp(p^2-16)+D(p^2-16)="

"=43p^3+104p^2+432p+1024"


"A+B+C=43"

"-4A+4B+D=104"

"16A+16B-16C=432\implies A+B-C=27"

"-64A+64B-16D=1024\implies -4A+4B-D=64"


"2C=16\implies C=8"

"2D=40\implies D=20"


"A+B=35"

"-4A+4B=84\implies B-A=21"

"B=28,A=7"


Then:

"Y(p)=7\cdot\frac{1}{p+4}+28\cdot\frac{1}{p-4}+\frac{8p+20}{p^2+16}"

"\frac{1}{p+4}\to e^{-4t}"

"\frac{1}{p-4}\to e^{4t}"

"\frac{p}{p^2+16}\to cos4t"

"\frac{1}{p^2+16}\to \frac{1}{4}sin4t"


Answer:

"y(t)=7e^{-4t}+28e^{4t}+8cos4t+5sin4t"